OpenStudy (anonymous):

please help PLEASE 0=-16t^2+96t-140

OpenStudy (anonymous):

factor and please show steps

jhonyy9 (jhonyy9):

first you can factor it out -2 ,yes ?

jhonyy9 (jhonyy9):

so than what will result ?

OpenStudy (anonymous):

i need to know what happens to the 24

jhonyy9 (jhonyy9):

24 ? from what you have got it ?

OpenStudy (aravindg):

\[\huge t=\frac{7}{2}\;\;\; or\;\;\; t=\frac{5}{2}\]

OpenStudy (anonymous):

0= -4(4t^2-24t+35)

OpenStudy (anonymous):

the 4 disappears

OpenStudy (anonymous):

divide both sides by 4, 0 divided by 4 =0

OpenStudy (anonymous):

yes i know but what is the next step ?

OpenStudy (anonymous):

0= -4(4t^2-24t+35)

OpenStudy (anonymous):

4t^2-24t+35= 0

OpenStudy (anonymous):

ok but the 24 how do u factor that?

OpenStudy (aravindg):

try this way \[\Large 16\times(t-3)^2 -4=0\]

OpenStudy (anonymous):

im confused

OpenStudy (aravindg):

where?

OpenStudy (anonymous):

at^2+bt+c=0 <-- compare this formula with urs, substitute the letters with the numbers and use this formula : t= (-b+-squareroot(b^2 - 4ac)) / 2a

OpenStudy (anonymous):

see when its factored out it turns into 0= -4(2t-5)(2t-7) so where does the 24 go? how does the 24 get factored?

OpenStudy (anonymous):

u dont have to factor if the equation is too complicatted, so just use the formula

OpenStudy (anonymous):

here check this tutorial out all the explainations are there http://tutorial.math.lamar.edu/Classes/Alg/SolveQuadraticEqnsI.aspx

OpenStudy (anonymous):

the quadratic formula always works, till now you'd probably finish it !! \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a }\]

OpenStudy (anonymous):

-16t^2+64t-28

OpenStudy (anonymous):

okay first you need to simplify it a bit. Just see that all the coefficients are divisible by 4 you will get -4t^2+24t-35 now you can see that there are no factors in whole numbers available for this equation. now you need to use quadratic formula \[-b \pm \sqrt{b^2 -4ac} /2a\]

OpenStudy (anonymous):

so i dont have to factor i can just use the quadratic formula?

OpenStudy (anonymous):

of course.

OpenStudy (anonymous):

simply it is (2t-7)(2t-5)=0 so t=7/2 or t=5/2